# PhysDreamer: Physics-Based Interaction with 3D Objects via Video Generation

### v1

@@ 1 Introduction
Realistic object interactions play a pivotal role in creating immersive virtual experiences. Recent advances in 3D vision have enabled the capture and creation of high-quality static 3D assets [50, 36], and some methods even extend to 4D assets [47, 59, 46], generating unconditioned dynamics. However, these approaches cannot handle interactive action-conditioned dynamics, which is crucial for synthesizing realistic object responses to novel interactions, such as external forces or agent manipulations.
@@
Action-conditioned dynamics differs from unconditional and text-conditioned dynamics in that it requires perceiving the physical material properties of objects and grounding the 3D motion prediction on these properties, such as object stiffness. Yet, estimating physical material properties is a challenging task due to the lack of ground-truth data, as measuring these properties for real objects is highly difficult. Real-life objects often exhibit complex, spatially-varying material properties, making the estimation problem even more challenging.
@@
Figure 1: (Left) Leveraging and distilling prior knowledge of dynamics from a pre-trained video generation model, we estimate a physical material field for the static 3D object. (Right) The physical material field allows synthesizing interactive 3D dynamics under arbitrary forces. We show rendered sequences from two viewpoints, with red arrows indicating force directions. Please see the videos on our project website for better visualization.
@@
In this work, we focus on synthesizing interactive 3D dynamics. Our goal is to transform static 3D objects into interactive ones that can respond to novel interactions in a physically plausible manner. To address this challenge, we propose PhysDreamer, a physics-based approach that leverages the object dynamics priors learned by video generation models. The key idea behind PhysDreamer is to distill these priors to estimate the physical material properties of static 3D objects. We hypothesize that video generation models, trained on large amounts of video data, implicitly capture the relationship between object appearance and dynamics. By leveraging this learned prior knowledge, PhysDreamer can infer the physical material properties that drive the dynamic behavior of objects, even in the absence of ground-truth material data (Fig. 1).
@@
PhysDreamer represents 3D objects using 3D Gaussians [36], models the physical material field with a neural field [71], and simulates 3D dynamics using the differentiable Material Point Method (MPM) [28]. The differentiable simulation and rendering allow for direct optimization of the physical material field and initial velocity field by matching pixel space observations. We focus on elastic dynamics and showcase PhysDreamer through diverse real examples, such as flowers, plants, a beanie hat, and a telephone cord.
We evaluate the realism of the synthesized interactive motion through a user study, comparing PhysDreamer to state-of-the-art methods. The results demonstrate that our approach significantly outperforms existing techniques in terms of motion realism, validating the effectiveness of leveraging video generation priors for estimating physical material properties and synthesizing interactive 3D dynamics.
@@ 2.1 Dynamic 3D reconstruction
Dynamic 3D reconstruction methods aim to reconstruct a representation of a dynamic scene from inputs such as depth scans[13, 41], RGBD videos[51], or monocular or multi-view videos [69, 52, 57, 53, 75, 1, 47, 40, 74, 45, 66]. This task is especially challenging in the monocular setting with slow-moving cameras and fast-moving scenes [20]. Novel scene representations are a major driver of recent progress. One prominent approach is to augment a canonical Neural Radiance Fields (NeRF) with a deformation field [57]. This approach can be further improved by incorporating flow supervision [66, 23] or as-rigid-as-possible or volume preserving regularization terms [52, 53]. Time-modulated NeRFs [43, 20, 19, 7] offer a simpler alternative representation. Due to its Lagrangian nature, 3D Gaussian Splatting[36] is readily adaptable to the task of efficient dynamic scene reconstruction [47, 40, 74, 72, 31, 17]. Data-driven prior, such as from monocular depth models [77, 39] and image diffusion models [67], can also be used to reduce the inherent ambiguity in dynamic reconstruction from monocular videos.
@@ 2.3 Interactive motion generation
For 3D assets, physics-based approaches enable synthesizing motions under any physical interactions. Virtual Elastic Objects[10] jointly reconstructs the geometry, appearances, and physical parameters of elastic objects in a multiview capture setup with compressed air system, allowing realistic animations under any new physical interactions. PAC-NeRF [42] and PhysGaussian [18] integrate physics-based simulations with NeRF and 3D Gaussians, respectively, to generate physically plausible motions. We use the same physics-based approach to generate realistic interactions, but a novel ingredient of our work is to distill the material parameters of the object from pre-trained video generation models.
@@ 2.4 Video generation models
Recent progress in video generation is driven by the development of larger autoregressive [65, 27, 68, 38] and diffusion models [60, 26, 5, 22, 4, 3, 6, 24]. These models, trained on increasingly large datasets, continue to advance the quality and realism of generated video content. The state-of-the-art approach [6] can generate minute-long videos with realistic motions and viewpoint consistency. However, current video generation models cannot support physics-based control, like interactions with objects through external forces. Our approach leverages strong motion prior from a video generation model, and as the field of video generation continues to advance, we expect our method to yield even better results.
@@ 3 Problem formulation
Therefore, our problem formulation focuses on estimating the spatially varying Young’s modulus field $E(\bm{x})$ for the 3D object. To allow particle simulation, we query a particle’s Young’s modulus by $E_{p}=E(\bm{x}_{p})$. As for other physical properties, the mass for a particle $m_{p}$ can be pre-computed as the product of a constant density ($\rho$) and particle volume $V_{p}$. The particle volume can be estimated [70] by dividing a background cell’s volume by the number of particles that cell contains. As for the Poisson’s ratio $\nu_{p}$, we found that it has negligible impact on object motion in our preliminary experiments, and so we assume a homogeneous constant Poisson’s ratio.
@@ 4.1 Preliminaries
Material Point Method (MPM). We use the Moving Least Squares Material Point Method (MLS-MPM) [28] to solve the above governing equation. MPM is a hybrid Eulerian-Langrangian method widely adopted for simulating dynamics for a wide range of materials, such as solid, fluid, sand, and cloth [58, 37, 62, 32]. MPM offers several advantages, such as easy GPU parallelization [29], handling of topology changes, and the availability of well-documented open-source implementations [30, 49, 48, 70].
@@
$\bm{C}_{p}^{t+1}=\frac{4}{(\Delta x)^{2}}\sum_{i}N(\bm{x}_{i}-\bm{x}_{p}^{t})\bm{v}_{i}^{t}(\bm{x}_{i}-\bm{x}_{p}^{t})^{T},\quad\bm{F}_{p}^{t+1}=(\bm{I}+\Delta t\bm{C}_{p}^{t+1})\bm{F}_{p}^{t}.$ (6)
@@ 4.2 Estimating physical properties
Using MLS-MPM [28, 70] as our physics simulator and the Fixed Corotated hyper-elastic material model for the 3D objects. the simulation process for a single sub-step is formalized as:
@@ 4.3 Accelerating simulation with subsampling
Specifically, we apply K-Means clustering to create a set of driving particles $\{\mathcal{Q}_{q}\}_{q=1}^{Q}$ at $t=0$, where each driving particle is represented by $\mathcal{Q}^{0}_{q}=\{\bm{x}^{0}_{q},\bm{v}^{0}_{q},\bm{F}^{0}_{q},\bm{C}^{0}_{q},E_{q},m_{q},\nu_{q},V_{q}\}$. The initial position of a driving particle $\bm{x}_{q}^{0}$ is computed as the mean of the position $\bm{x}_{p}$ of all cluster members. The number of the driving particles is much smaller than the number of 3D Gaussian particles, $Q\ll P$. We run simulations only on the driving particles. During rendering, we compute the position and rotation for each 3D Gaussian particle $\mathcal{G}_{p}$ by interpolating the driving particles. In particular, for each 3D Gaussian particle, we find its eight nearest driving particles at $t=0$, and we fit a rigid body transformation $\bm{T}$ between these eight driving particles at $t=0$ and at the current timestamp. This rigid body transformation $\bm{T}$ is applied to the initial position and rotation of the particle $\mathcal{G}_{p}$ to obtain its current position and rotation. We summarize our algorithm with pseudo-code in Appendix 0.C.
@@ 5.1 Setup
We recruited $100$ participants, each asked to judge all $8$ scenes, forming a total of $800$ 2AFC judgement samples for each baseline comparison. For each scene, we create $4$ sample video pairs and show participants a random one from the $4$ pairs. In particular, we create $4$ five-second motion sequences using PhysDreamer with randomized initial conditions (applying an external force to the foreground object or assigning an initial velocity to the object), and render videos from randomly picked viewpoints. For the baseline method, we apply the same initial conditions (for PhysGaussian only) and render videos from the same viewpoint as ours to form the video pairs. Please see Appendix 0.B for human study details, and Appendix 0.A for quantitative metrics for videos (e.g., Fréchet Video Distance[64]).
@@ 5.2 Implementation details
Neural material fields. We parameterize both our material field and initial velocity field using triplanes [55, 9] each followed by a three-layer MLP. The triplanes has spatial resolution of $8^{3}$ and $24^{3}$ for the material field and velocity field, respectively.
@@
Generating reference videos. We render a 3D object with its background from a viewpoint, and then we use Stable Video Diffusion [4] to animate this rendered image and generate fourteen video frames. We use a small motion bucket number [4] (e.g., 5 or 8) so that the generated video contains mostly object motion and little camera motion. We specifically use rendered images rather than Gaussians’ training images for the video generation, because we might not have access to training images (e.g., the 3D Gaussians are generated from text-to-3D models). Also, rendering images directly from 3D Gaussians simplifies later optimization. While not necessary, in our experiments, we choose a viewpoint from the original 3D Gaussian’s training set.
@@ 5.3 Results
Surprisingly, even when compared to real videos, $53.7\%$ 2AFC samples favored the motion realism of motion synthesized by PhysDreamer. However, viewers biased with prior knowledge, like the authors, can easily distinguish simulated videos from real videos, since PhysDreamer produces motion with lower frequencies in most experiments. This low-frequency motion pattern is evident in the space-time slice visualizations at Fig. 6. For a more detailed visual comparison, please see the supplementary materials.
@@ 5.4 Ablation: using multi-view reference videos
Figure 7: Comparison between single-view (top) and two-view (bottom) supervisions. The object (alocasia) exhibits self-occluding structures. We can use generated videos at two views to jointly optimize the material field. Single-view reference provides less effective supervision signals for materials of occluded parts, as shown in the blue boxes. In the space-time (X-t) slices, the vertical axis represents time, and the horizontal axis shows a spatial slice of the object.
For objects with self-occlusion, observing salient motion of all object parts from a single video is challenging (e.g., the alocasia scene where a leaf can occlude another leaf). We may alleviate this problem by rendering from multiple viewpoints to provide comprehensive coverage of the object. Here, we use multiple videos in the material estimation, jointly optimizing a video-agnostic, spatially-varying Young’s modulus for each particle along with video-specific initial velocities. From the comparison of the alocasia scene in Fig. 7, we can see that using multi-view reference videos (a front view and a back view) helps in such complex self-occluding objects: PhysDreamer benefits significantly from having supervision from two views, while using only a single view leads to artifacts. In our user study, $81.0\%$ 2AFC samples prefer PhysDreamer with two view supervision in visual quality and $86.0\%$ in motion realism.
@@ 6 Conclusion
In this work, we introduced PhysDreamer, a novel approach for synthesizing interactive 3D dynamics by endowing static 3D objects with physical material properties. Our method leverages the object dynamics priors learned by video generation models to estimate the spatially-varying material properties. The estimated material properties enable realistic physical interaction with the 3D objects. We demonstrated the effectiveness of PhysDreamer through experiments on a diverse set of elastic objects and a user study, showing that our approach significantly outperforms state-of-the-art methods in terms of motion realism. We believe that PhysDreamer takes a significant step towards creating more engaging and immersive virtual environments by enabling static 3D objects to respond realistically to novel interactions, opening up a wide range of applications from realistic simulations to interactive virtual experiences.
@@
This work is in part supported by the NSF PHY-2019786 (The NSF AI Institute for Artificial Intelligence and Fundamental Interactions), RI #2211258, #2338203, ONR MURI N00014-22-1-2740, Quanta Computer, and Samsung. We would like to thank Peter Yichen Chen, Zhengqi Li, Pingchuan Ma, Minghao Guo, Ge Yang, and Shai Avidan for help and insightful discussions.

### v2

@@ 1 Introduction
Realistic object interactions play a pivotal role in creating immersive virtual experiences. Recent advances in 3D vision have enabled the capture and creation of high-quality static 3D assets [53, 37], and some methods even extend to 4D assets [50, 62, 49], generating unconditioned dynamics. However, these methods fail to handle action-conditioned dynamics in response to new physical interactions, such as synthesizing the motion of a rose reacting to a breeze or a touch.
@@
The key challenge in synthesizing action-conditioned dynamics lies in understanding the physical material properties of objects. Yet, estimating these properties is a challenging task due to the lack of ground-truth data, as measuring these properties for real objects is highly difficult. Real-life objects often exhibit complex, spatially-varying material properties, making the estimation problem even more challenging. Despite the complexity of physical materials, humans can easily imagine how objects would react to external forces, such as the gentle sway of a rose. This ability to imagine object dynamics stems from our physical prior knowledge obtained from observing and interacting with the physical world. This motivates us to distill dynamics priors from video generation models that have been trained on vast, diverse video observations of the physical world.
@@
Figure 1: (Left) Leveraging and distilling dynamics priors from a pre-trained video generation model, we estimate a physical material field for the static 3D object. (Right) The physical material field allows synthesizing interactive 3D dynamics under arbitrary forces. We show rendered sequences from two viewpoints. Red arrows indicate force directions. Please see videos on our project website for better visualization.
@@
In this work, we focus on synthesizing interactive 3D dynamics. We propose PhysDreamer, a physics-based approach to transforming static 3D objects into interactive ones that can respond to novel interactions. The key idea behind PhysDreamer is to distill dynamics priors learned by video generation models to estimate the physical material properties of static 3D objects. We hypothesize that video generation models, trained on large amounts of video data, implicitly capture the relationship between object appearance and dynamics. By leveraging this learned prior knowledge, PhysDreamer can infer the physical material properties that drive the dynamic behavior of objects, even in the absence of ground-truth material data (Fig. 1).
@@
PhysDreamer represents 3D objects using 3D Gaussians [37], models the physical material field with a neural field [75], and simulates 3D dynamics using the differentiable Material Point Method (MPM) [35, 74]. The differentiable simulation and rendering allow for direct optimization of the physical material field and initial velocity field by matching pixel space observations. We focus on elastic dynamics and showcase PhysDreamer through diverse real examples, such as flowers, plants, a beanie hat, and a telephone cord. We evaluate the realism of the synthesized interactive motion through a user study, comparing PhysDreamer to state-of-the-art methods. The results demonstrate that our approach significantly outperforms existing techniques on motion realism and visual quality.
@@ 2.1 Dynamic 3D reconstruction
Dynamic 3D reconstruction methods aim to reconstruct a representation of a dynamic scene from inputs such as depth scans[14, 44], RGBD videos[54], or monocular or multi-view videos [73, 55, 60, 56, 79, 1, 50, 42, 7, 78, 48, 70]. This task is especially challenging in the monocular setting with slow-moving cameras and fast-moving scenes [21]. Novel scene representations are a major driver of recent progress. One prominent approach is to augment a canonical Neural Radiance Fields (NeRF) with a deformation field [60]. This approach can be further improved by incorporating flow supervision [70, 24] or as-rigid-as-possible or volume preserving regularization terms [55, 56]. Time-modulated NeRFs [46, 21, 20, 8] offer a simpler alternative representation. Due to its Lagrangian nature, 3D Gaussian Splatting[37] is readily adaptable to the task of efficient dynamic scene reconstruction [50, 42, 78, 76, 32, 18]. Data-driven prior, such as from monocular depth models [81, 41] and image diffusion models [71], can also be used to reduce the inherent ambiguity in dynamic reconstruction from monocular videos.
@@ 2.3 Interactive motion generation
For 3D assets, physics-based approaches enable synthesizing motions under any physical interactions. Virtual Elastic Objects[11] jointly reconstructs the geometry, appearances, and physical parameters of elastic objects in a multiview capture setup with compressed air system. PAC-NeRF [45], DANO [43], and PhysGaussian [19] integrate physics-based simulations with NeRF and 3D Gaussians to generate physically plausible motions. We use the same physics-based approach to generate realistic interactions, but a novel ingredient of our work is to distill the material parameters of the object from pre-trained video generation models.
@@ 2.4 Video generation models
Recent progress in video generation is driven by the development of larger autoregressive [69, 28, 72, 40] and diffusion models [63, 27, 5, 23, 4, 3, 6, 25]. These models, trained on increasingly large datasets, continue to advance the quality and realism of generated video content. The state-of-the-art approach [6] can generate minute-long videos with realistic motions and viewpoint consistency. However, current video generation models cannot support physics-based interactions with objects through external forces.
@@ 3 Problem formulation
Therefore, our problem formulation focuses on estimating the spatially varying Young’s modulus field $E(\bm{x})$ for the 3D object. To allow particle simulation, we query a particle’s Young’s modulus by $E_{p}=E(\bm{x}_{p})$. As for other physical properties, the mass for a particle $m_{p}$ can be pre-computed as the product of a constant density ($\rho$) and particle volume $V_{p}$. The particle volume can be estimated [74] by dividing a background cell’s volume by the number of particles that cell contains. As for the Poisson’s ratio $\nu_{p}$, we found that it has negligible impact on object motion in our preliminary experiments(see supplementary materials for details), and so we assume a homogeneous constant Poisson’s ratio.
@@ 4.1 Preliminaries
Material Point Method (MPM). We use the Material Point Method (MPM) [35, 74] to solve the above governing equation. MPM is a hybrid Eulerian-Langrangian method widely adopted for simulating dynamics for a wide range of materials, such as solid, fluid, sand, and cloth [61, 38, 66, 33]. MPM offers several advantages, such as easy GPU parallelization [30], handling of topology changes, and the availability of well-documented open-source implementations [31, 52, 51, 74].
@@
$\bm{C}_{p}^{t+1}=\frac{4}{(\Delta x)^{2}}\sum_{i}N(\bm{x}_{i}-\bm{x}_{p}^{t})\bm{v}_{i}^{t}(\bm{x}_{i}-\bm{x}_{p}^{t})^{T},\bm{F}_{p}^{t+1}=(\bm{I}+\Delta t\sum_{i}\bm{v}_{i}^{t}\nabla N(\bm{x}_{i}-\bm{x}_{p}^{t})^{T})\bm{F}_{p}^{t}.$ (6)
@@ 4.2 Estimating physical properties
Using MPM [35, 74] as our physics simulator and the Fixed Corotated hyper-elastic material model for the 3D objects, the simulation process for a single sub-step is formalized as:
@@ 4.3 Accelerating simulation with subsampling
Specifically, we apply K-Means clustering to create a set of driving particles $\{\mathcal{Q}_{q}\}_{q=1}^{Q}$ at $t=0$, where each driving particle is represented by $\mathcal{Q}^{0}_{q}=\{\bm{x}^{0}_{q},\bm{v}^{0}_{q},\bm{F}^{0}_{q},\bm{C}^{0}_{q},E_{q},m_{q},\nu_{q},V_{q}\}$. The initial position of a driving particle $\bm{x}_{q}^{0}$ is computed as the mean of the position $\bm{x}_{p}$ of all cluster members. The number of the driving particles is much smaller than the number of 3D Gaussian particles, $Q\ll P$. We run simulations only on the driving particles. During rendering, we compute the position and rotation for each 3D Gaussian particle $\mathcal{G}_{p}$ by interpolating the driving particles. In particular, for each 3D Gaussian particle, we find its eight nearest driving particles at $t=0$, and we fit a rigid body transformation $\bm{T}$ between these eight driving particles at $t=0$ and at the current timestamp. This rigid body transformation $\bm{T}$ is applied to the initial position and rotation of the particle $\mathcal{G}_{p}$ to obtain its current position and rotation. We summarize our algorithm with pseudo-code in supplementary materials.
@@ 5.1 Setup
We recruited $100$ participants, each asked to judge all $8$ scenes, forming a total of $800$ 2AFC judgement samples for each baseline comparison. For each scene, we create $4$ sample video pairs and show participants a random one from the $4$ pairs. In particular, we create $4$ five-second motion sequences using PhysDreamer with randomized initial conditions (applying an external force to the foreground object or assigning an initial velocity to the object), and render videos from randomly picked viewpoints. For the baseline method, we apply the same initial conditions (for PhysGaussian only) and render videos from the same viewpoint as ours to form the video pairs. Please see supplementary materials for human study details and quantitative metrics for videos (e.g., Fréchet Video Distance[68]).
@@ 5.2 Implementation details
Neural material fields. We represent both material field and initial velocity field using triplanes [58] each followed by a three-layer MLP. The triplanes have spatial resolutions of $8^{3}$ and $24^{3}$ for the material field and velocity field, respectively.
@@
Generating reference videos. We render a 3D object with its background from a viewpoint, and then we use Stable Video Diffusion [4] to animate this rendered image and generate fourteen video frames. We use a small motion bucket number [4] (e.g., 5 or 8) so that the generated video contains mostly object motion and little camera motion. We use rendered images for the video generation, so that our approach can also be used for generated scenes. Also, rendering images directly from 3D Gaussians simplifies later optimization.
@@ 5.3 Results
Compared to real videos, $53.7\%$ 2AFC samples favored the motion realism of ours results. Interestingly, under “Motion Realism”, 86% of the users indicated that the alocasia outputs were more realistic than real captures. This is surprising, as one would expect a 50% preference if the videos were indistinguishable. We offer a potential explanation: for thin geometries like alocasia leaves, the Material Point Method tends to produce lower-frequency and slower motions. This can be observed in the video and is evident in the space-time slice visualizations in Fig. 6. Humans are poor at judging the naturalness of motion and may be biased towards smoother and slower motions, as shown in prior studies [65, 39].
@@ 5.4 Ablation: using multi-view reference videos
Figure 7: Comparison between single-view (top) and two-view (bottom) supervisions. The object (alocasia) exhibits self-occluding structures. We can use generated videos at two views to jointly optimize the material field. In the space-time (X-t) slices, the vertical axis represents time, and the horizontal axis shows a spatial slice of the object.
For objects with self-occlusion, observing salient motion of all object parts from a single video is challenging (e.g., the alocasia scene where a leaf can occlude another leaf). We may alleviate this problem by rendering from multiple viewpoints to provide comprehensive coverage of the object. Here, we use multiple videos in the material estimation, jointly optimizing a video-agnostic, spatially-varying Young’s modulus for each particle along with video-specific initial velocities. From the comparison of the alocasia scene in Fig. 7, we can see that using multi-view reference videos (a front view and a back view) helps in such complex self-occluding objects: PhysDreamer benefits significantly from having supervision from two views, while using only a single view leads to artifacts. In our user study, $81.0\%$ 2AFC samples preferresults with two view supervision in visual quality and $86.0\%$ in motion realism.
@@ 6 Conclusion
In this work, we introduced PhysDreamer, a novel approach to synthesizing interactive 3D dynamics by endowing static 3D objects with physical material properties. Our method distills the object dynamics priors learned by video generation models to estimate the spatially-varying material properties. We showcased dynamics interaction with a diverse set of elastic objects by PhysDreamer. We believe that PhysDreamer takes a significant step towards creating more engaging and immersive virtual environments, opening up a wide range of applications from realistic simulations to interactive virtual experiences.
@@
This work is in part supported by the NSF PHY-2019786 (The NSF AI Institute for Artificial Intelligence and Fundamental Interactions, http://iaifi.org/), NSF CIF 1955864 (Occlusion and Directional Resolution in Computational Imaging), RI #2211258, #2338203, ONR MURI N00014-22-1-2740, Quanta Computer, Samsung, and United States Air Force Artificial Intelligence Accelerator and was accomplished under Cooperative Agreement Number FA8750-192-1000. We would like to thank Peter Yichen Chen, Zhengqi Li, Pingchuan Ma, Minghao Guo, Ge Yang, and Shai Avidan for help and insightful discussions.